If u multiply a number by 6 and then subtract negative ten, the difference is negative 38. What is the number?
Answer by ankor@dixie-net.com(22740)

If u multiply a number by 6 and then subtract negative ten, the difference is negative 38.
What is the number?
:
6x - (-10) = -38
minus a minus is a plus
6x + 10 = -38
6x = -38 - 10
6x = -48
x = -48/6
x = -8
:
:
Confirm this in the original equation
6(-8) - (-10) = -38
-48 + 10 = -38